The income elasticity is +2 and income increases by 20%.sales were 5000 units, what woill they be now

1
Expert's answer
2021-12-15T11:48:41-0500

"e_d=\frac{\Delta D/D}{\Delta I/I}\\2=\frac{\Delta D/5000}{0.2}\\0.4=\frac{\Delta D}{5000}\\\Delta D=5000*0.4=2000"

Therefore, the sales will now be:

"5000+2000=7000" units